ln(y+4)=ln(y+1)+ln5; I tried working each side out individually on my calculator but the answers do not agree with my any of my 4 choices. I came up with 2.639 for left side and 4.007 for right side
Answer by stanbon(75887)

ln(y+4) = ln(y+1)+ln5
----
ln(y+4) = ln[5(y+1)]
----
y+4 = 5(y+1)
y+4 = 5y+5
4y = -1
y = -1/4
====================
Cheers,
Stan H.